OpenStudy (anonymous):

If a basketball player has a vertical leap of 20 inches what is his hangtime? Use hangtime function of V=48T^2

OpenStudy (anonymous):

so I am trying ot use the equation of 20=48T^2

OpenStudy (anonymous):

pretty much

OpenStudy (anonymous):

Plug in the given "vertical leap" of 20 inches and solve for t: v=48t^2 20=48t^2 20/48 = t^2 0.645 seconds = t